OpenStudy (anonymous):

Use the Rational Root Theorem to list all possible rational roots of 3x3 + x2 – 15x – 5 = 0. Then find any actual rational roots. Find the roots of 10x4 + x3 + 7x2 + x – 3 = 0 A polynomial equation with rational coefficients has the roots Sq rt of 7 and –5i. Find two additional roots Find a third-degree polynomial equation with integer coefficients that has the roots 8 and 3i.

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

@satellite73

OpenStudy (jennychan12):

uhhhh, why is this in physics? u should put this in the math group. u'll get more help. not sure what the RRT is x^2(x+3)-5(x-1) = 0 so that it's (x^2-5)(x+3)(x-1) = 0 so that means that x = sqrt(5), -3, or 1. Find a third-degree polynomial equation with integer coefficients that has the roots 8 and 3i. (x-8)(x^2+3) sorry i can't help you with the rest... :'(

OpenStudy (anonymous):

ohh snap i just realized that.... i thought it was math